物理化学特論 2 単位 (選択)

Advanced Physical Chemistry  魚崎泰弘 · 教授/ 環境創生工学専攻化学機能創生コース物質合成化学講座，鈴木良尚 · 講師/ 環境創生工学専攻化学機能創生コース物質機能化学講座

【授業目的】溶液中で起こる様々な物理化学的現象，特に溶媒和，結晶成長というトピックスを通して，物理化学の基礎的な原理と実際の現象との結びつきについて理解を深める．

【授業概要】 (溶媒和) 液体中で化学反応を研究することが多い．それらの反応で，溶媒和が大きく関与するが，定量的に理解することは困難な場合が多い．そのため，統計力学に基づく「溶媒和熱力学」を導入して溶媒和を理論的に解釈する手法について講述する． (結晶成長) 結晶成長学は，材料設計や構造解析に対し重要な役割を果たしている．その結晶成長学を通して，平衡状態の熱力学，反応速度論，表面界面の化学などについてより深く学ぶ．

【授業形式】講義【キーワード】溶媒和，結晶成長【先行科目】『基礎物理化学』 (1.0) 【関連科目】『物質機能化学特論』 (0.5) 【履修要件】学部教育における物理化学を理解していること．【履修上の注意】授業を受ける際には， 2 時間の授業時間毎に 2 時間の予習と 2 時

間の復讐をしたうえで授業を受けることが，授業の理解と単位取得のために必要である．

【到達目標】 1. 溶媒和現象の基礎を理解する

2. 結晶成長の基礎を理解する【授業計画】

1. (溶媒和) 統計力学と熱力学

2. (溶媒和) 化学ポテンシャル

3. (溶媒和) 溶媒和熱力学

4. (溶媒和) 溶媒和エネルギー

5. (溶媒和) イオン溶媒和

6. (溶媒和) 超臨界流体中の固体の溶解度

7. (溶媒和) 超臨界流体中の溶媒和

8. (結晶成長) 相変化の熱力学

9. (結晶成長) 結晶の誕生

10. (結晶成長) 理想的成長

11. (結晶成長) 表面構造とラフニング

12. (結晶成長) 表面 kinetics 13. タンパク質の結晶化

14. タンパク質の結晶構造解析

15. コロイド結晶

【成績評価基準】授業中に課すレポート 100%で評価する．

【教科書】講義中に適宜紹介する

【参考書】齋藤幸夫著結晶成長 (裳華房フィジックスライブラリー)

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216869 【対象学生】開講コース学生のみ履修可能【連絡先】

⇒ 魚崎 (化 510, 088-656-7417, uosaki@chem.tokushima-u.ac.jp) ���� (オフィスアワー: 月曜日 17:00-18:00)

Advanced Physical Chemistry 2 units (selection)

Yasuhiro Uosaki · Professor / Synthetic and Polymer Chemistry, Chemical Science and Technology, Earth and Life Environmental Engineering

Yoshihisa Suzuki · Associate Professor / Physicochermistry and Material Science, Chemical Science and Technology, Earth and Life Environmental Engineering

Target⟩ The main goal of this class is to understand the relationship between the principle of physical chemistry and real phenomena. Solvation, crystal growth are the main topics of this class.

Outline⟩ (Solvation phenomena) Many chemical reactions are studied in liquids. Although solvation process plays a key role in these reactions, it is hard to understand the solvation quantitatively. In this lecture, ”solvation thermodynamics” based on statistical mechanics is introduced and the methods to understand the solvation theoretically are explained. (Crystal Growth) Concepts and technology of crystal growth play an important role in design of materials and structure analysis. The main goal of this class is to provide a solid introduction to the fundamentals of crystal growth that can be used to describe various phenomena involving equilibrium conditions, rate processes, surface or interface properties, etc.

Style⟩ Lecture �������⟩ solvation, crystal growth Fundamental Lecture⟩ “Basic Physical Chemistry”(1.0) Relational Lecture⟩ “Advanced Topics in Materials Science”(0.5) Requirement⟩ Students are required to have a good understanding of undergraduate-

level physical chemistry and related subjects. Notice⟩ 授業を受ける際には， 2 時間の授業時間毎に 2 時間の予習と 2 時間の復

讐をしたうえで授業を受けることが，授業の理解と単位取得のために必要である．

Goal⟩

1. To understand the fundamentals of solvation phenomena

2. To understand the fundamentals of crystal growth

Schedule⟩

1. Statistical mechanics and thermodynamics

2. Chemical potential

3. Solvation thermodynamics

4. Solvation energy

5. Ion solvation

6. Solubilities of solids in supercritical fluids 7. Solvation in supercritical fluids

8. Thermodynamics of phase transition

9. Nucleation

10. Ideal growth rate of crystal

11. Surface structure and roughning transition

12. Surface kinetics

13. Protein crystallization

14. Protein crystallography

15. Colloidal crystals

Evaluation Criteria⟩ Assignments count 100%.

Textbook⟩ To be announced in the class

Reference⟩ Yukio Saito, Statistical Physics of Crystal Growth, World Scientific,

Singapore, 1996

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216869

Student⟩ Able to be taken by only specified class(es)

Contact⟩

⇒ Uosaki (G510, +81-88-656-7417, uosaki@chem.tokushima-u.ac.jp) ���� (Office Hour: Monday (17:00-18:00))